A bag contains four $5 bills, five $10 bills and three $20 bills. two bills

are selected without the first selection being replaced find P($5 then

$5)


1
Expert's answer
2021-06-29T08:03:22-0400

Solution:

Number of $5 bills = 4

Number of $10 bills = 5

Number of $20 bills = 3

Total bills = 4+5+3 = 12

P($5 then $5) = P($5) x P($5)

"=\dfrac{4}{12} \times \dfrac{3}{11} = \dfrac{1}{11}"